OpenStudy (anonymous):

Find the slope of a line parallel to the line through the given points. C(0, 3), D(4, 6)

OpenStudy (anonymous):

is it 3/4?

OpenStudy (zepp):

\(\large \frac{Delta~y}{Delta~x}=\frac{6-3}{4-0}=\frac{3}{4}\) Yes :)

OpenStudy (anonymous):

but for a line parallel to this?

Parth (parthkohli):

If a line is parallel to these points, then the slope is same.

OpenStudy (lgbasallote):

parallel means same slope ;)

OpenStudy (anonymous):

-.-

Parth (parthkohli):

So the slope will be 3/4. :D

OpenStudy (anonymous):

omg.. i feel so stoupanfoicn iohlcneihmocih

Parth (parthkohli):

@zepp \Delta

OpenStudy (anonymous):

thanks guise-.-

OpenStudy (lgbasallote):

@seashell i feel ya brah *brofist*

OpenStudy (zepp):

\(\Delta \) :O omg

OpenStudy (anonymous):

im not a man :O

OpenStudy (lgbasallote):

:O ommigosh!! everything i've lived for is a lie o.O don't tell me you're not a seashell either?!!!

OpenStudy (anonymous):

-.- sadly im a seashell .-.

OpenStudy (zepp):

\(\huge \Omega \omega \Pi \pi\) :OOO

OpenStudy (anonymous):

dude omgomgogmogmgmmm^^^

Parth (parthkohli):

\(\phi \Phi \)

OpenStudy (anonymous):

^ :OOOOO

OpenStudy (anonymous):

Find the slope of a line perpendicular to a line through the given points. E(5, 7), F(3 , 1)

OpenStudy (anonymous):

how do i do this^

OpenStudy (lgbasallote):

get the slope of the two points first...you know how to do that?

OpenStudy (anonymous):

:OOOOOOOOOOOOOOO no

Parth (parthkohli):

And then find the negative of the slope's reciprocal.

OpenStudy (lgbasallote):

\[\huge slope = \frac{y_2 - y_1}{x_2 - x_1}\] do you know now?

OpenStudy (anonymous):

oh yahh -.-

OpenStudy (anonymous):

is it 1/3

OpenStudy (lgbasallote):

yes!!!